OpenStudy (anonymous):

Jake earns $7.50 per hour working at a local car wash. The function, ƒ(x) = 7.50x, relates the amount Jake earns to the number of hours he works. Write the inverse of this relation.

OpenStudy (anonymous):

@freckles pls help me this

OpenStudy (sparrow2):

f(x)=x/7.5

OpenStudy (anonymous):

Thanks yeah i got the same answer. i just want to make sure it.

OpenStudy (sparrow2):

u r welcome

OpenStudy (anonymous):

Write the inverse function for the function, ƒ(x) = x + 4. Then, find the value of ƒ -1(4). Type your answers in the box. can you help with that @sparrow2

OpenStudy (sparrow2):

y=x-4

OpenStudy (sparrow2):

f(4)=4-4=0

OpenStudy (anonymous):

Thanks @sparrow2